### v1

# On the Rudin-Blass ordering of measures

Abstract.
We study the Rudin-Blass (and the Rudin-Keisler) ordering on the finite additive measures on $\omega$. We propose a generalization of the notion of Q-point and selective ultrafilter to measures: Q-measures and selective measures. We show some symmetries between Q-points and Q-measures but also we show where those symmetries break up. In particular we present an example of a measure which is minimal in the sense of Rudin-Blass but which is not a Q-measure.
Key words and phrases:
P-points, P-measures, Additive Property, Q-points, Q-measures, Selective measures, Ruding-Blass ordering, Rudin-Keisler ordering
2020 Mathematics Subject Classification:
Primary: 03E05. Secondary: 03E35, 03E75, 28E15, 28A33, 46E27.
The third author was was supported by the Grant Agency of Charles University (project no. 219025) and partially by the Czech Academy of Sciences CAS (RVO 67985840).
Introduction
By a measure $\mu$ on $\omega$ we mean a finitely additive set function of values in $[0,1]$ and such that $\mu(\omega)=1$. Every ultrafilter $\mathcal{U}$ on the natural numbers can be seen as an example of a measure, by assigning $1$ to elements of $\mathcal{U}$, and 00 to the elements of the dual ideal. Some measures, the non-atomic ones, are far from being ultrafilters. Perhaps the most natural example of such measure is the following:
$\mu(A)=\lim_{n\to\mathcal{U}}\frac{|A\cap n|}{n},$
where $\mathcal{U}$ is an ultrafilter on $\omega$.
Presumably, at the beginning of the history of ultrafilters, the non-atomic measures on $\omega$, particularly the invariant ones, were studied as intensively as the atomic ones (i.e. the ultrafilters), see e.g. [Cho69b], [Fai72], [Cho69a]. At some point, the theory of ultrafilters flourished and expanded in many different directions, whereas the non-atomic measures were studied with a relatively modest intensity. These were investigated by Fremlin and Talagrand in [FT79], van Douwen in [vD92], and other people (see more references below).
The structure of ultrafilters on $\omega$ can be investigated in different contexts. One of the most fruitful approaches is through Rudin-Keisler and Rudin-Blass orderings (see [Laf96] and [Bla73]). Loosely speaking, an ultrafilter $\mathcal{U}$ is smaller (in Rudin-Blass or Rudin-Keisler order) than $\mathcal{V}$ if $\mathcal{U}$ contains no more information than $\mathcal{V}$.
Although Rudin-Blass and Rudin-Keisler orderings can be easily generalized for measures, it seems that this generalization was considered for the first time very recently, in [BNCMM25], where it was used to study measures in the Silver extension. The authors proved that consistently there is a measure that does not have an ultrafilter Rudin-Keisler below it. This suggests that the order structure of measures does not necessarily come down to the ultrafilter case.
The most natural question concerning the RB and RK orderings is the one concerning their minimal elements. For ultrafilters, the Rudin-Blass minimal ultrafilters are exactly the Q-points and the Rudin-Keisler ones are exactly the selective ultrafilters. An ultrafilter $\mathcal{U}$ is a Q-point if every partition of $\omega$ into finite sets has a selector in $\mathcal{U}$. An ultrafilter $\mathcal{U}$ is selective if every partition into sets from the dual ideal has a selector in $\mathcal{U}$. These notions have natural generalizations to the realm of measures that, to the best of our knowledge, have not been studied yet (with a very recent exception of [AMCAS25] which we will mention later).
We say that a measure $\mu$ on $\omega$ is a Q-measure if every partition of $\omega$ into finite sets has a selector of full measure. A measure $\mu$ is selective if for every partition $(A_{n})$ of $\omega$ there is a selector $S$ of the biggest possible measure (i.e. $\mu(S)=1-\sum_{n}\mu(A_{n})$).
We managed to prove that there are strong symmetries between the theory of Q-measures, selective measures and P-measures (generalizations of P-points, see below). In particular, we proved the following results which are generalizations of the fact well known for ultrafilters:
  • under CH there are non-atomic selective measures,
  • a measure is selective if and only if it is simultaneously a P-measure and a Q-measure,
  • every Q-measure (selective measure) is minimal in Rudin-Blass (Rudin-Keisler, respectively) ordering,
  • the generic forcing for adding a non-atomic measure adds a selective measure.
There are also substantial differences. The main difference we found is that the property of being Q-measure is not equivalent to minimality in the Rudin-Blass ordering.
Theorem 5.5.
Assuming CH, there is a Rudin-Blass minimal measure which is not a Q-measure (and there is an Rudin-Keisler minimal measure which is not selective).
This measure has interesting properties. It is far from being a Q-measure: it is shift-invariant and Lebesgue measurable (as a function from the Cantor set to the reals). Q-measures are Rudin Blass minimal because its filters of measure 1 sets are sufficiently close to Q-points. The reason for the minimality of the above measure is more delicate and follows from its ’fractal’ structure.
We also found an interesting dichotomy concerning a weakening of a Q-measure: A measure is a Q+-measure if for each partition into finite sets there is a selector of positive measure (this is a natural counterpart of Q+ filters). We found the following result:
Theorem 5.3.
A Rudin-Blass minimal measure is either a Q+-measure or it is Rudin-Blass equivalent to a shift-invariant measure.
So, a measure which is Rudin-Blass minimal but is not Q+ has to be (equivalent to a measure) symmetric with respect to the natural semigroup operation.
The Q-points and selective ultrafilters are often put in one line with another type of ultrafilter: a P-point (e.g. selective ultrafilters are exactly those which are Q-points and P-points simultaneously). This notion has been generalized to the measure context in 1984, by Mekler (see [Mek84]). We call this generalization a P-measure (originally it was called measures with AP property). It turned out to be an interesting notion which attracted attention of several mathematician (see [BFPRN01], [Gre19], [Kun22], [BNS23]).
It is easy that the existence of P-points implies the existence of a non-atomic P-measure. It was a longstanding open problem (until recently) if the existence of P-measures implies the existence of P-points. It turns out that the answer is negative. Interestingly, the situation for Q-measures is quite the opposite. We have summarized this duality in the following table.
Type $\exists$ ult. $\implies$ $\exists$ nonatomic measure $\exists$ measure $\implies$ $\exists$ ult.
P yes ([Mek84]) no ([BNCMM25])
Q no (Corollary 4.28) yes (Proposition 4.12)
selective no (Corollary 4.28) no (Theorem 4.25)
While we were working on this article we were informed by Antonio Avilés about the research in [AMCAS25], in which the authors also defined Q-measures (although they were interested more in Q+-measures). Interestingly, the authors of [AMCAS25] were motivated by a certain problem in Banach space theory. We overview some of their results in Section 4.2.
In Section 1 we introduce basic definitions and notation. In Section 2 we present some ways of constructing measures on $\omega$ which will be used later on in the article. In Section 3 we define Rudin-Blass and Rudin-Keisler orderings on measures. In Section 4 we introduce the special types of measures: P-measures, Q-measures, and selective measures, we discuss their basic properties and their existence. Section 5 is devoted mostly to the construction of a RB-minimal measure which is shift-invariant.
1. Preliminaries
By a measure on $\omega$ we always mean a finitely additive probability measure. We will also assume that it vanishes on points, i.e. $\mu(F)=0$ for every finite $F\subseteq\omega$. We will call such measures free but we will often assume this property without mentioning it. If $\mathcal{U}$ is an ultrafilter, then the function $\delta_{\mathcal{U}}$ assigning $1$ to elements of $\mathcal{U}$ and 00 to the rest, is a measure. Sometimes we will just treat ultrafilters as measures. Notice that an ultrafilter is free if and only if it is free as a measure.
If $A,B\subseteq\omega$, then $A\subseteq^{*}B$ denotes that $|A\setminus B|<\omega$. Usually we will not distinguish sets which are equal modulo finite sets (and so, formally, we will work in $\mathcal{P}(\omega)/fin$). A pseudo-intersection of a family $\mathcal{A}\subseteq\mathcal{P}(\omega)$ is a set $X\subseteq\omega$ such that $X\subseteq^{*}A$ for each $A\in\mathcal{A}$.
Every (finitely additive) measure on a Boolean algebra $\mathbb{A}$ can be uniquely extended to a $\sigma$-additive measure on the Stone space of $\mathbb{A}$ (see [Ple24, Corollary 2.8]). Also, every $\sigma$-additive measure defined on a compact zerodimensional space $X$ induces uniquely a (finitely additive) measure on the Boolean algebra of clopen subsets of $X$. In particular, (free) measures on $\omega$ can be extended to $\sigma$-additive measures on $\beta\omega$ ($\beta\omega\setminus\omega$) and every measure on $\beta\omega$ ($\beta\omega\setminus\omega$) induces a (free) measure on $\omega$.
A measure $\mu$ on a Boolean algebra $\mathbb{A}$ is non-atomic if for every $\varepsilon>0$ there is a partition $(A_{n})_{n<N}$ of the unity such that $\mu(A_{n})<\varepsilon$ for every $n<N$. Notice that a measure $\mu$ is non-atomic if and only if its extension to $\sigma$-additive measure on the Stone space vanishes on points.
A measure $\mu$ on $\omega$ is shift-invariant if $\mu(A+1)=\mu(A)$ for each $A\subseteq\omega$, where $A+1=\{n+1\colon n\in A\}$.
By $\mathbb{B}$ we will denote the measure algebra $\mathbb{B}=\mathrm{Bor}(2^{\omega})/_{\lambda=0}$, where $\lambda$ is the Lebesgue measure on $2^{\omega}$. The measure $\lambda$ can be naturally transported to $\mathbb{B}$, and the resulting measure is also denoted by $\lambda$.
[…]
Let $\mathcal{M}$ be the set of all bounded signed measures on $\omega$. This set can be endowed with the norm $\|\mu\|=\sup\{\mu(A)-\mu(\omega\setminus A):\,A\subseteq\omega\}$, forming a real Banach space. It can be shown that $\mathcal{M}\cong(\ell_{\infty})^{*}$. More precisely, the following transformations are isomorphism (given $\mu\in\mathcal{M}$ and $\varphi\in(\ell_{\infty})^{*}$):
$\varphi_{\mu}(x)=\int x\,\mathrm{d}\mu,\quad\quad\mu_{\varphi}(A)=\varphi(\chi_{A}).$
More details on this construction can be found in [BRBR83].
The other natural topology on $\mathcal{M}$ is the weak topology. Recall that the sets of the form
[…]
for $\varepsilon>0$, $\mu\in\mathcal{M}$ and a bounded function $f\colon\omega\to\omega$, form subbasis for this topology. In fact, we may assume that $f$ is a characteristic function so the sets
$K(\varepsilon,\mu,A)=\{\nu\colon|\mu(A)-\nu(A)|<\varepsilon\}$
for $\varepsilon>0$, $\mu\in\mathcal{M}$ and $A\subseteq\omega$, also form a subbasis for the same topology.
To finish this section, we present the following basic property of measures.
Proposition 1.1.
Let $\mu$ be a measure and let $F$ be the filter of sets of measure $1$, then $F\subseteq\mathcal{P}(\omega)$ is non-meager.
Proof.
Notice that the quotient $\mathcal{P}(\omega)/F^{*}$ is c.c.c, where $F^{*}$ is the dual ideal. So, if $F$ is meager, then, by Talagrand’s theorem (see [Tal80] or [BJ95, Theorem 4.1.2]) there must be an interval partition $\omega=\bigcup_{n\in\omega}P_{n}$ such that, for all $B\in F$ and almost all $n\in\omega$, we have that $B\cap P_{n}\neq\emptyset$. So, if $\mathcal{A}$ is an uncountable almost disjoint family, then $\{\bigcup_{i\in A}P_{i}:A\in\mathcal{A}\}$ is an uncountable antichain in $\mathcal{P}(\omega)/F^{*}$. ∎
Below, we present several ways of producing concrete non-atomic measures on $\omega$.
2. Constructions of measures
As we have mentioned in the introduction, every ultrafilter can be seen as a measure. For an ultrafilter $\mathcal{U}$ define the Dirac delta $\delta_{\mathcal{U}}(A)$ to be $1$ if $A\in\mathcal{A}$, otherwise 00. Sometimes in what follow we will just identify ultrafilters with its Dirac deltas. In this article we are more interested in non-atomic measures. By a result of Rudin [Rud57], every compact non-scattered space carries a non-atomic measure. Since $\beta\omega$ is one of such spaces, there is a non-atomic measure on $\omega$.
2.1. Extensions of density.
Recall that the asymptotic density is defined as
$d(A)=\lim_{n\to\infty}\frac{|A\cap n|}{n}$
provided the limit exists. The asymptotic density behaves very much like a measure, it is just undefined for many subsets of $\omega$. The simplest way of extending it to a measure is by using an ultrafilter limit.
Let $\mathcal{U}$ be an ultrafilter on $\omega$. Define the function $d_{\mathcal{U}}\colon\mathcal{P}(\omega)\to[0,1]$, an ultrafilter density, by
$d_{\mathcal{U}}(A)=\lim_{n\to\mathcal{U}}\frac{|A\cap n|}{n}.$
It is straightforward to check that $d_{\mathcal{U}}$ is a measure on $\omega$ which vanishes on points (provided $\mathcal{U}$ is free) and which is non-atomic. It extends density, i.e. $d_{\mathcal{U}}(A)=d(A)$ when the latter is defined.
This approach can be generalized. For a measure $\mu$ on $\omega$ we can define
[…]
Notice that for an ultrafilter $\mathcal{U}$ we have $d_{\delta_{\mathcal{U}}}=d_{\mathcal{U}}$. So, we may define a measure $d_{\mathcal{U}}$, then $d_{d_{\mathcal{U}}}$ and so on. However, some measures defined as above cannot be achieved in this process:
Proposition 2.1.
There is a measure $\mu$ such that $\mu=d_{\mu}$.
Proof.
Consider the mapping $d\colon\mathcal{M}\to\mathcal{M}$ defined by $d(\mu)=d_{\mu}$. Notice that this mapping is continuous. Indeed, fix $\mu$, $\varepsilon>0$ and $A\subseteq\omega$ and let $f$ be the sequence $(|A\cap n|/n)$. Take $\mu^{\prime}\in d^{-1}[K(\varepsilon,\mu,A)]$. Then
[…]
Hence, $K(\varepsilon-\delta,\mu^{\prime},f)\subseteq d^{-1}[K(\varepsilon,\mu,A)]$ and so $d$ is continuous.
Since $\mathcal{M}$ is convex and compact (by Banach-Alaoglu theorem), we can use Schauder-Tychonoff fixed point theorem to conclude that there is $\mu$ such that $\mu=d(\mu)=d_{\mu}$. ∎
We finish this subsection by recalling the following well-known fact.
Proposition 2.2.
If $\mu$ extends the density, then $\mu$, as a function from the Cantor set to $[0,1]$, is Lebesgue measurable.
Proof.
By Borel’s normal number theorem $\lambda(\{A\colon d(A)=1/2\})=1$. ∎
2.2. Transfinite constructions a’la Sikorski
The easiest way to obtain certain types of ultrafilters is to perform a construction under Continuum Hypothesis using transfinite induction of length $\omega_{1}$. As we will see, a similar situation holds for measures. In fact, we will not construct measures, but rather construct an embedding of $\mathcal{P}(\omega)$ into a measure algebras (with measure $\lambda$), as if $\phi$ is such embedding, then $\mu(A)=\lambda(\varphi(A))$ is a measure on $\omega$.
Let $\mathbb{C}$ be a complete Boolean algebra and let $\mathbb{A}\subseteq\mathcal{P}(\omega)$ be a Boolean algebra. For a Boolean homomorphism $\phi\colon\mathbb{A}\to\mathbb{C}$ define
$\phi^{*}(X)=\bigwedge\{\phi(A)\colon A\in\mathbb{A},\,\,X\subseteq A\}$
and
$\phi_{*}(X)=\bigvee\{\phi(A)\colon A\in\mathbb{A},\,\,A\subseteq X\}$
for $X\in\mathcal{P}(\omega)$. By $\mathrm{alg}(\mathcal{A})$ we denote the Boolean algebra generated by $\mathcal{A}$ and let $\mathbb{A}(X)=\mathrm{alg}(\mathbb{A}\cup\{X\})$.
The following lemma will let us create the measure ‘step by step’, by transfinite induction. The proof can be found in [Sik69].
Lemma 2.3.
([Sik69]). Let $\mathbb{A}\subseteq\mathcal{P}(\omega)$ be a Boolean algebra that is closed under finite modifications, let $\mathbb{C}$ be a complete Boolean algebra and $\phi\colon\mathbb{A}\to\mathbb{C}$ be a homomorphism. If $X\subseteq\omega$ and $C\in\mathbb{C}$ is such that
$\phi_{*}(X)\subseteq C\subseteq\phi^{*}(X),$
then $\phi$ can be extended to a homomorphism $\phi^{\prime}\colon\mathbb{A}(X)\to\mathbb{C}$ such that $\phi^{\prime}(X)=C$.
We will show now the simplest application of the above for the purpose of constructing a measure on $\omega$. Recall that the measure algebra $\mathbb{B}$ is complete.
Proposition 2.4.
There is a Boolean endomorphism $\phi\colon\mathcal{P}(\omega)\to\mathbb{B}$.
The usual ZFC proof of Proposition 2.4 is to start with an independent family $\mathcal{A}$ of subsets of $\omega$, of size $\mathfrak{c}$. Define a bijection $f\colon\mathcal{A}\to\mathbb{B}$ which, by the fact that $\mathcal{A}$ is independent, can be automatically extended to a homomorphism $\phi\colon\mathrm{alg}(\mathcal{A})\to\mathbb{B}$ and then extend $\phi$ to a homomorphism $\phi\colon\mathcal{P}(\omega)\to\mathbb{B}$.
However, under $\mathsf{CH}$, there is more flexibility: Start with a trivial homomorphism $\phi_{0}\colon\{0,\omega\}\to\mathbb{B}$ and enumerate $\mathbb{B}=(B_{\alpha})_{\alpha<\omega_{1}}$. Then, for $\alpha<\omega_{1}$, use Lemma 2.3 to extend $\phi_{\alpha}$ to $\phi_{\alpha+1}$ in such a way that $B_{\alpha}$ is in the range of $\phi_{\alpha+1}$. To see that this is possible let =$\mathcal{L}=\{X\subseteq\omega\colon\phi(X)\subseteq B_{\alpha}\}$, =$\mathcal{R}=\{X\subseteq\omega\colon B_{\alpha}\subseteq\phi(X)\}$. Both $\mathcal{L}$ and $\mathcal{R}$ are countable and for each $X\in\mathcal{L}$, $Y\in\mathcal{R}$ we have $X\subseteq Y$. So, there is $Z\subseteq\omega$ such that $X\subseteq^{*}Z\subseteq^{*}Y$ for each $X\in\mathcal{L}$, $Y\in\mathcal{R}$, since $(\omega,\omega)$-gaps do not exists (see [Sch93]). At a limit step $\gamma$, for $\phi_{\gamma}$ we take the union of all the homomorphisms defined so far.
The advantage of the proof using transfinite induction is that it can be easily modified to prove existence of homomorphisms with some additional properties. For example, the following was proved in [Kun76]:
Proposition 2.5.
(CH) There is a Boolean endomorphism $\phi\colon\mathcal{P}(\omega)\to\mathbb{B}$ whose kernel is a P-ideal.
To prove it one has to add another ’promise’ in the construction described above: for every countable $\mathcal{A}$ such that $\phi(A)=1$ for each $A\in\mathcal{A}$, there is a pseudo-intersection $P$ of $\mathcal{A}$ such that $\phi(P)=1$. It is not hard to see (can be found in [Kun76]) that this promise can be satisfied along the transfinite construction described above and that the resulting homomorphism $\phi$ will have the desired property. We will use this scheme, modifying some of these promises.
2.3. Generic measure.
First, recall a standard fact about extending measures on Boolean algebras (which can be seen as a corollary of Lemma 2.3, see also e.g. [Pla76]).
Lemma 2.6.
Suppose that $\mu$ is a measure defined on a Boolean algebra $\mathbb{A}\subseteq\mathcal{P}(\omega)$ closed under finite modifications, and that $X\subseteq\omega$. Let $r\in[0,1]$ be such that $\mu_{*}(X)\leq r\leq\mu^{*}(X)$. Then $\mu$ can be extended to a measure $\nu$ on $\mathbb{A}(X)$ such that $\nu(X)=r$.
We present a natural example of a generic measure. Recall that the forcing $\mathcal{P}(\omega)/\mathrm{Fin}$ (with the reverse inclusion ordering) adds generically an ultrafilter. The following is its measure counterpart.
Example 2.7.
We consider the family of all (free) partial measures on $\omega$:
=$\mathbb{P}=\{\mu\colon\mu\mbox{ is a free measure on a countable Boolean algebra }\mathbb{A}\subseteq\mathcal{P}(\omega)\}$
with the ordering given by $\mu\leq\nu$ if $\mu\supseteq\nu$ (so, if the domain of $\nu$ is a subalgebra of the domain of $\mu$ and the measures agree on its elements).
Then $\mathbb{P}$ is a forcing notion which is $\sigma$-closed: for every decreasing sequence $(\mu_{n})$, there is $\mu\in\mathbb{P}$ such that $\mu\leq\mu_{n}$ for each $n$. The forcing $\mathbb{P}$ adds generically a measure $\dot{\mu}$ on $\omega$ (as for every $X\subseteq\omega$ and a countable Boolean algebra $\mathbb{A}\subseteq\mathcal{P}(\omega)$ we can always extend $\mu$ to $\mathbb{A}(X)$, using Lemma 2.6).
[…]
2.4. Solovay measures.
In [BNS23] the authors dust off an old construction of Solovay (see [Sol71]) of measures in the random model. We will sketch the construction (for the details, see [BNS23]).
Fix an ultrafilter $\mathcal{U}$ on $\omega$. We will construct a $\mathbb{B}$-name for a measure on $\omega$ generated by $\mathcal{U}$.
Let $\dot{M}$ be an $\mathbb{B}$-name for a subset of $\omega$. Define the function $\mu_{\dot{M}}\colon\mathbb{B}\to[0,1]$ as follows:
$\mu_{\dot{M}}(B)=\lim_{k\to\mathcal{U}}\lambda\big{(}\llbracket k\in\dot{M}\rrbracket\wedge B\big{)}.$
[…]
Now let $\dot{r}_{\dot{M}}$ be the value of $f_{\dot{M}}$ on the generic real. Finally, let $\dot{\mu}_{\mathcal{U}}$ be a $\mathbb{B}$-name for a function $\mathcal{P}(\omega)\to[0,1]$ such that for every $\mathbb{B}$-name $\dot{M}$ for a subset of $\omega$ we have
$\Vdash_{\mathbb{M}_{\kappa}}\dot{\mu}_{\mathcal{U}}(\dot{M})=\dot{r}(\mu_{\dot{M}}).$
In [Sol71] it is proved that
$\Vdash_{\mathbb{B}}``\dot{\mu}_{\mathcal{U}}\mbox{ is a finitely additive probability measure on }\omega".$
[…]
We call $\dot{\mu}_{\mathcal{U}}$ evaluated by a generic the Solovay measure (associated to $\mathcal{U}$).
2.5. Transfinite constructions a’la Mokobodzki
Under CH (or, more generally, under the assumption that $\mathrm{cov}(\mathcal{M})=\mathfrak{c}$) there is a measure $\mu$ on $\omega$ which is universally measurable, i.e. it is such that it is $\rho$-measurable as a function $\mu\colon 2^{\omega}\to[0,1]$ for every Borel measure $\rho$ on the Cantor set. Such a measure is called a medial measure or Mokobodzki measure and it was constructed by Mokobodzki ([Mey73]) and independently by Christensen ([Chr74]) (see also [God16]).
We will present the idea of the construction and in Section 5 we will use it to define a peculiar example of a measure on $\omega$.
It will be convenient for us to change our setting and to use functionals rather than measures (see Preliminaries). We will recursively construct sequences of functionals on $\ell^{\infty}$ with finite support. Note that any uniformly bounded sequence of functionals $\langle c_{n}\rangle_{n}$ can be treated as a linear function from $\ell^{\infty}$ to $\ell^{\infty}$ by putting
$c(x)(n)=c_{n}(x).$
Define $e_{n}:\ell^{\infty}\to\mathbb{R}$ by $e_{n}(x)=x(n)$.
Definition 2.8.
Let $\langle c_{n}\rangle_{n},\langle d_{n}\rangle_{n}$ be two sequences of bounded functionals on $\ell^{\infty}$. We say that $\langle c_{n}\rangle_{n}\preceq\langle d_{n}\rangle_{n}$ whenever for all $n$ we have $c_{n}\in conv(d_{m}:m\geqslant n)$. Furthermore, let $\langle c_{n}\rangle_{n}\preceq^{*}\langle d_{n}\rangle_{n}$ whenever for all but finitely many $n$ we have $c_{n}\in conv(d_{m}:m\geqslant n)$. Both $\preceq$ and $\preceq^{*}$ are preorders.
Definition 2.9.
A functional $c:\ell^{\infty}\to\mathbb{R}$ has finite support if it is a convex combination of $e_{n}$’s, i.e. it is of the form
$c=\sum_{n\in F}a_{n}e_{n},$
where each $a_{n}$ is non-negative and $\sum_{n\in F}a_{n}=1$. We will call the set $F$ the support of $c$.
Observe that, if $\langle c_{n}\rangle_{n}\preceq\langle e_{n}\rangle_{n}$, then each $c_{n}$ has finite support. Additionally, for any $r\in\mathbb{R}$, $c_{n}(\vec{r})=r$, where $\vec{r}$ is the sequence constantly equal $r$. The following fact is the hearth of the construction.
Fact 2.10.
If $\langle c_{n}\rangle_{n}\preceq\langle d_{n}\rangle_{n}$ and $\lim_{n}d_{n}(x)$ exists, then $\lim_{n}c_{n}(x)=\lim_{n}d_{n}(x)$.
The Mokobodzki construction is performed under CH, by first defining a sequence $\langle\langle c_{n}^{\alpha}\rangle_{n}:\alpha\in\omega_{1}\rangle$ which is $\preceq$-increasing. Along the way the limits of $c^{\alpha}(x)$ exist for more and more $x\in\ell^{\infty}$. Finally we define a functional by $\varphi(x)=\lim_{n}c_{n}^{\alpha}(x)$ for the smallest $\alpha$ where such limit exists.
As an illustration of the above, we can perform the simplest construction of this form. Assume CH and enumerate =$\ell^{\infty}=\{x^{\alpha}\colon\alpha<\omega_{1}\}$. Start with $c_{n}^{0}=e_{n}$ and suppose we have defined $\langle c_{n}^{\alpha}\rangle_{n}$. There is a sequence $k_{n}$ such that $\lim c^{\alpha}_{k_{n}}(x_{n})$ does exist. Define $c_{n}^{\alpha+1}=c_{k_{n}}^{\alpha}$ for each $n$. If $\alpha$ is a limit ordinal, then fix a cofinal sequence $(\alpha_{n})_{n}$ and find $(n_{k})$ such that $\lim c^{\alpha_{n}}_{k_{n}}(x^{\alpha}_{n})$ does exist and proceed as in the successor case. Fix an increasing sequence $\langle\alpha_{n}\rangle_{n}$, cofinal in $\alpha$. Clearly $\langle c^{\alpha}_{n}\rangle_{n}$ is an $\preceq^{*}$-increasing sequence and for each $x\in\ell^{\infty}$ there is $\alpha$ such that $\lim c^{\alpha}_{n}(x)$ does exist. Finally, we obtain a functional, $\varphi(x)=\lim_{n}c_{n}^{\alpha}(x)$ for the smallest $\alpha$ where such limit exists (see 2.10).
The above construction is in fact nothing else but a limit along an ultrafilter (when we were picking subsequences, we actually chose subsets in an ultrafilter). So, the measure obtained in this way is just the Dirac delta of an ultrafilter. If instead we start with $c_{n}^{0}=(e_{1}+\dots e_{n})/n$ and proceed in the above way, then we would get a measure of the form $\mu(A)=\lim_{n\to\mathcal{U}}|A\cap n|/n$.
The construction of Mokobodzki imposes more conditions on the sequences $\langle c^{\alpha}_{n}\rangle_{n}$ to obtain universal measurability of the resulting functional. In Section 5 we will add yet another conditions to obtain a measure which is minimal with respect to Rudin-Blass ordering and which is not a Q+-measure.
3. Rudin-Blass and Rudin-Keisler orderings on measures
We start with a definition inspired by the well-known notion of Rudin-Blass ordering on filters (see [LZ98]). By $f[\mu]$ we mean the measure $\nu$ such that $\nu(A)=\mu(f^{-1}[A])$ for each $A$.
Definition 3.1.
Let $\mu,\nu$ be measures on $\omega$. We say that $\nu$ is Rudin–Blass reducible to $\mu$ (denoted by $\nu\leq_{RB}\mu$), if there is a finite-to-1 function $f\colon\omega\to\omega$ such that $f[\mu]=\nu$.
Definition 3.2.
We say that measure $\mu$ is nearly ultra if there exists an ultrafilter $\mathcal{U}$ such that $\delta_{\mathcal{U}}\leq_{RB}\mu$.
It is known that every ultrafilter extension of the asymptotic density is nearly ultra. Also, consistently (under Filter Dichotomy) every measure is nearly ultra. However, consistently there is a measure which is not nearly ultra (for the proofs of all these statements, see [BNCMM25]).
We say that a measure $\mu$ is Rudin-Blass minimal (RB-minimal, in short) if and only if for any $\nu$ such that $\nu\leq_{RB}\mu$ we have $\mu\leq_{RB}\nu$. It is known that an ultrafilter is a Q-point if and only if it is minimal in the Rudin–Blass ordering for ultrafilters.
Definition 3.3.
Let $\mu,\nu$ be measures on $\omega$. We say that $\nu$ is Rudin–Keisler reducible to $\mu$ (denoted by $\nu\leq_{RK}\mu$) if there is a function $f\colon\omega\to\omega$ such that $f[\mu]=\mu$.
So, Rudin-Blass ordering is just Rudin-Keisler ordering with the ’finite-to-one’ requirement. Again, the Rudin-Keisler ordering was investigated for ultrafilters and it is known that being a selective ultrafilter is equivalent to being minimal in this ordering (see [Bla70]).
The structure of the both Rudin-Blass and Rudin-Keisler orderings on ultrafilters has been extensively studied (see e.g. [RV19], [KR18]).
Proposition 3.4.
Suppose that $f\colon\omega\to\omega$ and $\mu$ is such that $f[\mu]=\mu$. Then for every $N\subseteq\omega$ we have $\mu(f[N])\geq\mu(N)$.
Proof.
Since $N\subseteq f^{-1}[f[N]]$ and $\mu(f^{-1}[A])=\mu(A)$ for every $A$, the conclusion follows. ∎
Proposition 3.5.
Suppose that $\mu$ is a measures on $\omega$. Then, there is a finite-to-one function $g\colon\omega\to\omega$ such that $g[\mu]=g[d_{\mu}]$ and, consequently, there is $\rho\leq_{RB}\mu,d_{\mu}$.
Proof.
Let $f\colon\omega\to\omega$ be a function such that $\lim_{n}f(n+1)/f(n)=\infty$ and let $I_{n}=\big{[}f(n),f(n+1)\big{)}$ for each $n$.
By the fact that $\{A\colon d_{\mu}(A)=0\}$ is non-meager, we may find an infinite $N\subseteq\omega$ such that $d_{\mu}\big{(}\bigcup_{k\in N}I_{k}\big{)}=0$. Let $(F_{n})$ be a sequence of intervals such that $N=\bigcup_{n}F_{n}$ and $\max F_{n}+1<\min F_{n+1}$ for every $n$. Finally, define $J_{n}=\big{[}f(\min F_{n}),f(\min F_{n+1})\big{)}$ and $K_{n}=\big{[}f(\max F_{n}),f(\min F_{n+1})\big{)}$.
Claim. For each $A\subseteq\omega$ we have $\mu\big{(}\bigcup_{A}J_{i}\big{)}=d_{\mu}\big{(}\bigcup_{A}J_{i}\big{)}$.
Fix $A\subseteq\omega$. First, notice that we have $\mu(\bigcup_{A}J_{i})=\mu(\bigcup_{A}K_{i})$ and so also $d_{\mu}(\bigcup_{A}J_{i})=d_{\mu}(\bigcup_{A}K_{i})$. Denote $B=\bigcup_{A}K_{i}$. For each $\varepsilon>0$ there is $m$ such that for every $n>m$ we have
$\left|\frac{|B\cap n|}{n}-\chi_{B}(n)\right|<\varepsilon.$
By the fact that $\mu$ vanishes on points, it means that
[…]
To conclude the proof of the proposition, define $g\colon\omega\to\omega$ by $g(k)=i$ if $k\in J_{i}$. ∎
Corollary 3.6.
For every ultrafilter $\mathcal{U}$, there is an ultrafilter $\mathcal{V}$ such that $\mathcal{V}\leq_{RB}d_{\mathcal{U}},\mathcal{U}$.
4. Special kinds of measures
4.1. P-measures
In this section we present the notion of P-measure, the natural generalization of P-points to the measure context. It was studied by several people, starting with Mekler [Mek84] (as AP measures, or measures with the additive property), Blass, Frankiewicz, Plebanek and Ryll-Nardzewski [BFPRN01], Grebik [Gre19], Kunisada [Kun22].
First, recall the definition of a P-filter.
Definition 4.1.
We say that a filter $\mathcal{F}$ is a P-filter if for each family $(A_{n})$ such that $A_{n}\in\mathcal{F}$ for all $n\in\omega$, there is a set $X\in\mathcal{F}$ which is a pseudo-intersection of $(A_{n})$. We say that a filter $\mathcal{U}$ is a P-point if it is an ultrafilter and a P-filter.
The following definition appeared in [Mek84]:
Definition 4.2.
A measure $\mu$ is a P-measure if for every descending sequence $(A_{n})$ there is a pseudo-intersection $P$ of $(A_{n})$ such that $\mu(P)=\lim\mu(A_{n})$.
The above definition can be rewritten equivalently in the language of partitions (which will be slightly more convenient for our purposes). We say that a set $S$ is a pseudo-selector of a partition $(P_{n})$ of $\omega$ if $S\cap P_{n}$ is finite for each $n$.
Proposition 4.3.
A measure $\mu$ is a P-measure if and only if for every partition $(P_{n})$ there is a pseudo-selector $S$ of $(P_{n})$ such that
$\mu(S)=1-\sum_{n}\mu(A_{n}).$
Notice that an ultrafilter $\mathcal{U}$ is a P-point if and only if $\delta_{\mathcal{U}}$ is a P-measure. The standard example of a P-measure which is non-atomic is this given in Section 2.1:
Theorem 4.4.
[Mek84] Let $\mathcal{U}$ be a P-point. Then the ultrafilter density
$d_{\mathcal{U}}(A)=\lim_{n\to\mathcal{U}}\frac{|A\cap n|}{n}$
is a non-atomic P-measure on $\omega$.
In [BNCMM25] the authors gave a fundamentally different example of a P-measure. Namely, they proved that the measure induced by the homomorphism from Proposition 2.5 is a P-measure. It follows from the fact that the filter of measure $1$ sets is a P-filter and that the Lebesgue measure $\lambda$ is $\sigma$-additive on $\mathbb{B}$.
The main result of [Mek84] was that consistently there are no P-measures. Mekler asked if the existence of P-measures imply the existence of P-points. The problem was solved recently in [BNCMM25]:
Theorem 4.5.
[BNCMM25] Consistently, there is a P-measure but there are no P-points.
In [Mek84] Mekler introduced another definition (under the name of AP0 measures):
Definition 4.6.
A measure $\mu$ is a P0-measure if for every descending sequence $(A_{n})$ of measure 1, there is a pseudo-intersection $P$ of $(A_{n})$ such that $\mu(P)=\lim\mu(A_{n})$.
Notice that an ultrafilter $\mathcal{U}$ is a P-point if and only if $\delta_{\mathcal{U}}$ is a P-measure, if and only if $\delta_{\mathcal{U}}$ is a P0-measure. So, in fact P0-measures are also possible generalizations of P-points to the context of measures. Notice that $\mu$ is a P0-measure if and only if the filter $\{F\colon\mu(F)=1\}$ is a P-filter. So, the property of being a P0-measure is more a property of the filter of measure 1 sets.
Proposition 4.7.
There exists a P0-measure that is not a P-measure.
Proof.
Notice that if $\mu$ is a P-measure, then it is $\sigma$-additive on the Boolean algebra $\mathcal{P}(\omega)_{/\mu=0}$. Consider the Cohen algebra $\mathbb{C}=\text{Bor}(2^{\omega})_{/\text{Meager}}$. It is a complete Boolean algebra. Moreover there exists a strictly positive measure on $\mathbb{C}$ and this measure is not $\sigma$-additive. Denote it by $\lambda^{\prime}$ (see [BND13] for more details). Hence we can create a measure $\mu^{\prime}$ on $\omega$ by transfinite induction which is a P0-measure and it is not $\sigma$-additive on $\mathcal{P}(\omega)_{/\mu=0}$, thus it is not a P-measure. ∎
4.2. Q-measures
First, recall the definition of a Q-point.
Definition 4.8.
We say that an ultrafilter $\mathcal{U}$ is a Q-point if for any $(A_{n})$, a partition of $\omega$ into finite sets, there exists $S\in\mathcal{U}$, such that $|S\cap A_{n}|\leq 1$.
The natural extension of this definition for measures seems to be the following.
Definition 4.9.
We say that $\mu$ is a Q-measure if for any $(P_{n})$, a partition of $\omega$ into finite sets, there exists $S$, such that $|S\cap P_{n}|\leq 1$ and $\mu(S)=1$.
As with P0-measures, this definition is rather about the filter of measure $1$ sets.
Proposition 4.10.
A measure $\mu$ is a Q-measure if and only if $\{F\subseteq\omega\colon\mu(F)=1\}$ is a Q-filter.
On the other hand, the selector $S=\{s_{0},s_{1},\dots\}$ from Definition 4.9 has the following property of preserving the measure: let $N\subseteq\omega$ be arbitrary. Denote $A=\bigcup_{n\in N}A_{n}$. We have
$\mu\big{(}\{s_{n}:n\in N\}\big{)}=\mu(S\cap A)=\mu(A)=\mu\Big{(}\bigcup_{n\in N}A_{n}\Big{)}.$
We will call such selector a measure preserving selector.
We will collect some easy observations.
Proposition 4.11.
If $\mathcal{U}$ is a Q-point, then $\delta_{\mathcal{U}}$ is a Q-measure.
Contrary to the case of P-measures, the existence of Q-measures (even non-atomic) implies the existence of Q-points.
Proposition 4.12.
The existence of Q-measures imply the existence of Q-points. Therefore, consistently there are no Q-measures.
Proof.
The first part follows from Proposition 4.10 and the obvious fact that every ultrafilter extending a Q-filter is a Q-point. The second follow from the classical results on models without Q-points (see e.g. [Mil80]). ∎
In fact, we can prove a slightly stronger fact. Recall that a measure is a Q+-measure if for each partition into finite sets we may find a selector of positive measure. The following theorem was proved independently in [AMCAS25].
Theorem 4.13.
Consistently, there are no $Q^{+}$-measures.
Proof.
Assume Filter Dichotomy and suppose that there is a $Q^{+}$-measure $\mu$. Let =$\mathcal{F}=\{F\colon\mu(F)=1\}$. Then, $\mathcal{F}$ is non-meager (Proposition 1.1). By Filter Dichotomy, there is an ultrafilter $\mathcal{U}\leq_{RB}\mathcal{F}$. We will show that $\mathcal{U}$ is a Q-point. Indeed, let $f$ be the finite-to-one function witnessing $\mathcal{U}\leq_{RB}\mathcal{F}$ and let $g\colon\omega\to\omega$ be any finite-to-one function. As $\mathcal{F}$ is a Q+-filter, there is a selector $S$ for $\big{(}(g\circ f)^{-1}(n)\big{)}_{n}$ such that $\mu(S)>0$. Then $S^{\prime}=f[S]$ is a selector for $(g^{-1}(n))_{n}$ and $S^{\prime}\in\mathcal{U}$. Otherwise $\mu(f^{-1}[S^{\prime}])=0$ which yields a contradiction as $S\subseteq f^{-1}[S^{\prime}]$. But Filter Dichotomy implies that there are no Q-points (see [Bla86, Corollary 15]). ∎
Of course, the inevitable question is if there is a model with a non-atomic Q-measure. Recall that there was a default way of getting a non-atomic P-measure from a P-point by extending the asymptotic density, see Theorem 4.4. However, in this way we will never get a Q-measure, in fact even Q+ measures cannot be shift-invariant (see Proposition 5.4).
Below we present an example of a non-atomic Q-measure, which will be obtained by the scheme described in Section 2.
Lemma 4.14.
Let $\mathbb{A}$ be a countable Boolean algebra on $\omega$ (closed under finite modifications), let $\mathbb{B}$ be a complete Boolean algebra, and let $(P_{n})$ be a partition of $\omega$ into finite sets. Let $\phi\colon\mathbb{A}\to\mathbb{B}$ be a homomorphism, such that $\phi(A)=\mathbf{0}$ for any finite $A\in\mathbb{A}$. There exists $S\subseteq\omega$, such that $|S\cap P_{n}|\leq 1$ for every $n\in\omega$ and
$\phi^{*}(S)=\mathbf{1}$
Therefore $\phi$ can be extended to a homomorphism $\phi^{\prime}\colon\mathbb{A}(S)\to\mathbb{B}$ such that
$\phi^{\prime}(S)=\mathbf{1}.$
Proof.
It suffices to show that there exists $S$, such that $|S\cap P_{n}|\leq 1$ for every $n\in\omega$ and for any $A\in\mathbb{A}$, if $S\subseteq\mathbb{A}$, then $\phi(A)=\mathbf{1}$.
Let $(A_{n})_{n\in\omega}$ be a sequence of all sets from $\phi^{-1}[\mathbb{B}\setminus\{\mathbf{1}\}]$. Let $f\colon\omega\to\omega$ be a function such that $x\in P_{f(x)}$. Now we will recursively construct $S$. We set
$s_{i}\in A_{i}^{c}\setminus\Big{(}\bigcup_{j<i}P_{f(s_{j})}\Big{)}.$
Note that it is possible as the set on the right-hand side is non-empty.
Finally set $S=\{s_{i}\colon i\in\omega\}$. It is easy to see that if $S\subseteq A$, then $\phi(A)=\mathbf{1}$. ∎
Theorem 4.15.
Assume $\mathsf{CH}$. There exists a non-atomic Q-measure.
Proof.
We proceed as in the schema presented in Section 2. Namely, we construct a homomorphism $\phi\colon\mathcal{P}(\omega)\to\mathbb{B}$ which fulfills the following conditions:
  • (1)
    for every $B\in\mathbb{B}$ there is $X\subseteq\omega$ such that $\phi(X)=B$,
  • (2)
    for every $(P_{n})_{n\in\omega}$, a partition of $\omega$ into finite sets, there exists a selector $S$ of this partition, such that $\phi(S)=\mathbf{1}$.
This way we can obtain a measure $\mu$ defined as $\mu(X)=\lambda(\phi(X))$ for every $X\subseteq\omega$. It follows from (1) that $\mu$ is non-atomic and from (2) that $\mu$ is a Q-measure.
Let $(\mathcal{P}_{\alpha})_{\alpha<\omega_{1}}$ be a sequence of all partitions of $\omega$ into finite sets. For the steps of the form $\gamma+2k$, $\gamma$ limit (or 00), $k\in\omega$, we can use Lemma 4.14 to find a selector $S$ of partition $\mathcal{P}_{\gamma+k}$ such that $\phi^{*}_{\gamma+2k}(S)=\mathbf{1}$ and set $\phi_{\gamma+2k+1}(S)=\mathbf{1}$. For the steps of the form $\gamma+2k+1$, we can take care of condition (1), in the way described in Section 2.2.
Then $\phi=\bigcup_{\alpha<\omega_{1}}\phi_{\alpha}$ satisfies (1) and (2). ∎
Recall that an ultrafilter is a Q-point if and only if it is RB-minimal. We will show that partially we have an analogous result for measures.
Proposition 4.16.
If $\mu$ is a Q-measure and $\nu\leq_{RB}\mu$, then $\mu\leq_{RB}\nu$.
Proof.
Let $f\colon\omega\to\omega$ be a finite-to-1 function witnessing $\nu\leq_{RB}\mu$. Denote $A=f[\omega]=\{a_{1},a_{2},\ldots\}$. Then $\nu(A)=\mu(f^{-1}[A])=\mu(\omega)=1$.
For $i\in\omega$ let $A_{i}=f^{-1}[\{a_{i}\}]$. Obviously, $(A_{i})$ is a partition of $\omega$ into finite sets and $\mu(A_{i})=\nu(\{a_{i}\})=0$ for every $i$.
Let $S=\{s_{i}\}_{i\in\omega}$ be a measure preserving selector of $(A_{i})$ promised by the fact that $\mu$ is a Q-measure. Define $g_{0}\colon A\to S$ by $g_{0}(a_{i})=s_{i}$. Clearly, $g_{0}$ is a bijection with $f_{\upharpoonright S}$ being its inverse and can be extended to a finite-to-1 function $g\colon\omega\to\omega$, such that $g[\omega]=S$.
Finally, we show that $g$ witnesses that $\mu\leq_{RB}\nu$. Let $X\subseteq\omega$. Then we have
$\nu(g^{-1}[X])$ $=\nu(g^{-1}[X\cap S])=\nu((g^{-1}[X\cap S]\cap A)\cup(g^{-1}[X\cap S]\cap A^{c}))$
$=\nu(g^{-1}[X\cap S]\cap A)+\nu(g^{-1}[X\cap S]\cap A^{c})$
$=\nu(g^{-1}[X\cap S]\cap A)=\nu(g_{0}^{-1}[X\cap S])$
$=\nu(f[X\cap S])=\mu(f^{-1}[f[X\cap S]])=\mu((X\cap S)\cap S)$
$=\mu(X).$
In the next section we will show that the above implication, in general, cannot be reversed.
Now, we overview the main results about Q-measures contained in [AMCAS25].
Theorem 4.17.
([AMCAS25]) There is no Q+-measure in the Laver’s model for no Q-points.
In fact the authors proved a more general result saying that no $\sigma$-bounded-cc ideal is rapid in the above model.
Also, the authors prove a result analogous to the well-known fact for ultrafilters saying that if $\mathfrak{d}=\mathrm{cov}(\mathcal{M})$, then every filter generated by $\mathfrak{d}$ many sets can be extended to a Q-point.
Theorem 4.18.
([AMCAS25]) The assumption $\mathfrak{d}=\mathrm{cov}(\mathcal{M})$ is equivalent to the statement: every atomless measure generated by less than $\mathfrak{d}$ many sets can be extended to a Q-measure.
4.3. Selective measures
First, recall the definition of selective ultrafilter.
Definition 4.19.
A filter $\mathcal{F}$ is selective if for every $(A_{n})$, a partition of $\omega$, such that $A_{n}^{c}\in\mathcal{F}$ for each $n$, there exists a selector $S$ of this partition, such that $S\in\mathcal{F}$.
Again, it seems that the natural generalization for measures is the following:
Definition 4.20.
A measure $\mu$ is selective if for every $(A_{n})$, a partition of $\omega$, there exists a selector $S$ of this partition, such that $\mu(S)=1-\sum_{n}\mu(A_{n})$.
Note that the selector $S$ from the definition above is a measure preserving selector.
It is a standard fact that the selective ultrafilters are exactly those which are P-points and Q-points. The analogous fact holds for measures.
Proposition 4.21.
A measure $\mu$ is selective if and only if it is a P-measure and a Q-measure.
Proof.
()($\Longrightarrow$)( ⟹ ). Follows directly from the definitions.
[…]
We may see that selective measures are RK-minimal. However, the proof is similar to the one given for Theorem 4.16, so we leave the next theorem without an explicit proof.
Definition 4.22.
A measure $\mu$ is 0-selective if for every $(A_{n})$, a partition of $\omega$ into sets of measure 00, there exists a selector $S$ of this partition, such that $\mu(S)=1$.
Theorem 4.23.
If $\mu$ is a 0-selective measure and $\nu\leq_{RK}\mu$, then $\mu\leq_{RK}\nu$.
It is not clear to us if the implication can be reversed.
Note that a selective measure can be constructed under $\mathsf{CH}$ in the similar way to Q-measure (see Theorem 4.15). Below we present some forcing constructions.
We consider the forcing $\mathbb{P}$ from Example 2.7:
=$\mathbb{P}=\{\mu\colon\mu\mbox{ is a free measure on a countable Boolean algebra }\mathbb{A}\subseteq\mathcal{P}(\omega)\}$
Proposition 4.24.
Let $\dot{\mu}$ be the generic measure added by $\mathbb{P}$. Then
$1\Vdash\dot{\mu}\mbox{ is selective.}$.
Proof.
We will use Proposition 4.21.
[…]
Theorem 4.25.
Consistently, there is a selective measure but there are no selective ultrafilters.
Proof.
We will work in the classical random model. Let $\mathcal{U}$ be a selective ultrafilter in the ground model and let $\dot{\mu}$ be the Solovay measure induced by $\mathcal{U}$ (see Section 2.4). We will work in $V[G]$. By [BNS23, Theorem 4.15] $\dot{\mu}$ is a P-measure. We have $\mathcal{U}\subseteq\{X\colon\dot{\mu}=1\}$ (in fact, by [BNS23, Theorem 5.6], we have even equality here). Since the random forcing is $\omega^{\omega}$-bounding, $\mathcal{U}$ is a Q-filter in $V[G]$ and so, by Proposition 4.10, $\dot{\mu}$ is a Q-measure. By Proposition 4.21 we have that $\dot{\mu}$ is a selective measure.
But there are no selective ultrafilters in the random model ([Kun76]). ∎
Theorem 4.26.
Suppose that there are infinitely many selective ultrafilters which are pairwise RK-incomparable. Then there is a non-atomic selective measure.
Proof.
Let $(\mathcal{U}_{n})$ be a family of pairwise RK-incomparable selective ultrafilters and let $\mathcal{U}$ be an ultrafilter. Let $\mu$ be a measure in the weak closure of the set of measures of the form
$\frac{1}{n}\big{(}\delta_{\mathcal{U}_{0}}+\dots+\delta_{\mathcal{U}_{n-1}}\big{)}.$
This is a non-atomic measure on $\omega$. To show that it is selective consider a function $f\colon\omega\to\omega$ and let $S_{n}$ be a selector of $(f^{-1}(n))_{n}$ appropriate for $\mathcal{U}_{n}$.
Claim. There is a family $(A_{n})$ of pairwise disjoint sets such that $f^{-1}[A_{n}]\in\mathcal{U}_{n}$ for each $n$.
Indeed, since $(\mathcal{U}_{n})$ are RK-incomparable, $(f[\mathcal{U}_{n}])_{n}$ are pairwise different. For every $n$, by the fact that $f[\mathcal{U}_{n}]$ is a P-point (and so it is not in a closure of any countable subset), there is $A_{n}\in f[\mathcal{U}_{n}]$ such that $A_{n}\notin f[\mathcal{U}_{k}]$ for $k\neq n$. The family $(A_{n})$ satisfies the desired property.
Let
$S=\bigcup_{n}\big{(}S_{n}\cap f^{-1}[A_{n}]\big{)}.$
The set $S$ is a selector of $(f^{-1}(n))_{n}$ and $\mu(S)=1$. ∎
Note that the assumption on non-comparability is crucial. Consider a selective ultrafilter $\mathcal{U}$ and for $n\in\omega$, define $\mathcal{U}_{n}$ as the ultrafilter generated by $\{n+A\colon A\in\mathcal{U}\}$. Then every measure in the weak closure of the set of measures of the form
$\frac{1}{n}\big{(}\delta_{\mathcal{U}_{0}}+\dots+\delta_{\mathcal{U}_{n-1}}\big{)}$
is shift-invariant and so it cannot be selective (see Proposition 5.4).
Proposition 4.27.
If there is an atomless Q-measure, then there are $2^{\mathfrak{c}}$ many coherence classes with Q-points.
Proof.
If $\mu$ is a non-atomic Q-measure, then =$\mathcal{F}=\{F\colon\mu(F)=1\}$ is a Q-filter. Since every closed infinite subset of $\beta\omega$ contains a copy of $\beta\omega$, there are $2^{\mathfrak{c}}$ extensions of $\mathcal{F}$ to ultrafilters. Clearly, all of them are Q-points. Now, notice that if Q-points are nearly coherent, then they are isomorphic (see e.g. [Bla73, Corollary 2]). But there are only $\mathfrak{c}$ many possible isomorphism classes. ∎
Corollary 4.28.
Consistently, there is a selective ultrafilter, but there are no non-atomic Q-measures.
Proof.
In [Mil24] Mildenberger presented a model in which there are exactly three coherence classes of ultrafilters on $\omega$ and there is a selective ultrafilter. By Proposition 4.27 there is no non-atomic Q-measure in this model. ∎
5. RB-minimal measures versus Q-measures
In this section we address the following question:
Problem 5.1.
Is every RB-minimal measure a Q-measure?
Of course, this is the case for ultrafilters. The main tool in the proof is the following fact (see [Bla73, Theorem 1]).
Proposition 5.2.
Suppose that $f\colon\omega\to\omega$ and an ultrafilter $\mathcal{U}$ is such that $f[\mathcal{U}]=\mathcal{U}$. Then the set of fix points of $f$ is in $\mathcal{U}$.
This fact cannot be directly generalized for measures: Consider $f(n)=n+1$ and a shift-invariant measure $\mu$. Then $f[\mu]=\mu$ but the set of fixed points is empty.
There is another possible way of generalizing Proposition 5.2 to the realm of measures. Again, suppose that $\mu$ is a measure and $f\colon\omega\to\omega$ is a finite-to-one function such that $f[\mu]=\mu$. We may define an equivalence relation on $\omega$:
(1) $k_{1}\sim k_{2}\mbox{ if there are }n_{1},n_{2}\mbox{ such that }f^{(n_{1})}(k_{1})=f^{(n_{2})}(k_{2}).$
Choose a transversal $S\subseteq\omega$ of equivalence classes of $\sim$. Define
$X=\bigcup\left\{f^{(n)}[S]\colon n>0\right\}.$
The set $X$ is an attractor of $f$, i.e. for each $k\in\omega$ there is $n$ such that $f^{(n)}(k)\in X$. It is also a selector for $\{f^{-1}[n]\colon n\in\omega\}$. It seems natural to conjecture that $\mu(X)=1$. That would be a natural generalization for Proposition 5.2. However this is not true:
Consider the measure from Example 2.1, i.e. $\mu$ such that
[…]
Let $f\colon\omega\to\omega$ be such that $f^{-1}[n]=[2^{n},2^{n+1})$. Then $f[\mu]=\mu$ since
[…]
But the equivalence relation (1) has only one equivalence class and so each attractor $X$ defined as above is finite and so $\mu(X)=0$.
The above approach reveals an interesting phenomenon: measures which are RB-minimal and which are not Q+-measures are RB-equivalent to a shift invariant measure.
Theorem 5.3.
Suppose $\mu$ is RB-minimal. If $\mu$ is not a Q+-measure, then there is a shift invariant $\nu\leq_{RB}\mu$.
Proof.
Let $f\colon\omega\to\omega$ be a finite-to-one function witnessing that $\mu$ is not Q+-measure. By minimality, we may assume that $f[\mu]=\mu$.
Let $S$ be a selector of the equivalence relation (1). If possible we choose an element $s$ with $f^{k}(s)=s$ for some $k$. Now let $\widehat{S}=\{f^{k}(n):n\in S,\,k\in\omega\}$. Notice that $\widehat{S}$ is a selector for $f$ and so $\mu(\widehat{S})=0$. We can arbitrarily change values on $\widehat{S}$ while preserving the properties of $f$.
Let $\widehat{f}$ be equal to $f$ on $\omega\setminus\widehat{S}$. Enumerate $\widehat{S}$ as $\{s_{n}:n\in\omega\}$ and let $\widehat{f}(s_{n+1})=s_{n}$ and $\widehat{f}(s_{0})=s_{0}$.
Notice, that $\widehat{f}$ is a finite-to one function with the following properties:
  • $\widehat{f}[\mu]=\mu$,
  • every selector for $\widehat{f}$ has measure 0,
  • for every $n\in\omega$ there is $k$ such that $\widehat{f}^{k}(n)=s_{0}$.
For each $n\in\omega$, let $k(n)$ denote the minimal $k$ as above.
For each $n$ the set $\{m:k(m)=n\}=k^{-1}[n]$ is finite and so $k$ is finite-to-one. We claim that $\nu=k[\mu]$ is shift-invariant.
Indeed, notice that $k^{-1}[n+1]=\widehat{f}^{-1}[k^{-1}[n]]$. Thus for every $A\subseteq\omega$ we have
$\nu(A+1)=\mu(k^{-1}[A+1])=\mu\big{(}\widehat{f}^{-1}[k^{-1}[A]]\big{)}=\mu(k^{-1}[A])=\nu(A).$
Compare the above theorem with the following proposition.
Proposition 5.4.
If $\mu$ is $Q^{+}$, then it is neither shift-invariant nor does it extend the density function.
Proof.
Let $I_{n}=[2^{n},2^{n+1})$ and let $S=\{s_{0}<s_{1}<\dots\}$ be a selector for $\langle I_{n}\rangle_{n}$ of positive measure. Notice that $|S\cap k|<\log_{2}(k)$ and so $S$ has density 0.
Assume that $\mu$ is shift-invariant. Let $S^{\prime}=\{s_{2k}\colon k\in\omega\}$ and assume (throwing out the first element of $S$ if necessary) that $\mu(S^{\prime})>0$. Take $n$ such, that $\mu(S^{\prime})>\frac{1}{n}$. Notice that the sets $S^{\prime},S^{\prime}+1,S^{\prime}+2,\cdots S^{\prime}+n$ are pairwise almost disjoint and of the same measure. Thus $\mu(\bigcup_{k=0}^{n}(S^{\prime}+k))>1$ which is a contradiction. ∎
5.1. Construction of a shift-invariant RB-minimal measure
We are going to prove the following theorem.
Theorem 5.5.
Under CH there is a shift-invariant RB-minimal measure. Consequently, under CH there is an RB-minimal measure that is not a Q+-measure.
We will prove this in a series of lemmas. We will use the notation of Section 2.5. In particular we will deal with sequences $\langle c_{n}\rangle_{n}$ of functionals on $\ell^{\infty}$ such that $\langle c_{n}\rangle_{n}\preceq\langle e_{n}\rangle_{n}$. Notice that it means that each $c_{n}$ is a convex combination of $e_{m}$’s. Consequently each $c_{n}$ has a finite support and $c_{n}(\chi_{A})=1$ iff $\mathrm{supp}(c_{n})\subseteq A$ and $c_{n}(\chi_{A})=0$ iff $\mathrm{supp}(c_{n})\cap A=\emptyset$.
Lemma 5.6.
For any sequence of functionals $\langle c_{n}\rangle_{n}\preceq\langle e_{n}\rangle_{n}$ and a finite-to-one function $f$, there is a subsequence $\langle c^{\prime}_{n}\rangle_{n}$ and a finite-to-one function $g$ having the property
$(\forall{x\in l_{\infty}})\;c^{\prime}_{n}(x\circ g\circ f)=x(n).$
Proof.
We will inductively construct $g$ in such a way that $g^{-1}[\{n\}]$ are going to be consecutive intervals. Let $g^{-1}[\{0\}]$ be long enough so that the support of $c_{0}$ is contained in $(g\circ f)^{-1}[\{0\}]$ and put $c_{0}^{\prime}=c_{0}$. For $n>0$ let $k_{n}=1+\max(g\circ f)^{-1}[\{0,1,\cdots n-1\}]$. Since $\langle c_{n}\rangle_{n}\preceq\langle e_{n}\rangle_{n}$, the support of $c_{k_{n}}$ is disjoint with $[0,k_{n})$ and so also with $(g\circ f)^{-1}[\{0,1,\cdots n-1\}]$. Let $g^{-1}[\{n\}]$ be long enough so that the support of $c_{k_{n}}$ is contained in $(g\circ f)^{-1}[\{n\}]$ and put $c^{\prime}_{n}=c_{k_{n}}$.
Note that, if $A\subseteq\omega$, then we have that $c^{\prime}_{n}(\chi_{A}\circ g\circ f)=c_{k_{n}}(\chi_{(g\circ f)^{-1}[A]})=\chi_{A}(n)$ since $\operatorname{supp}c_{k_{n}}\subseteq(g\circ f)^{-1}[\{n\}]$ and $\operatorname{supp}c_{k_{n}}\cap(g\circ f)^{-1}[\{m\}]=\emptyset$ for $m\neq n$. By linearity and continuity, this equality extends to all $x\in\ell^{\infty}$. ∎
For the following proofs, we will work with sequences of functionals. For simplicity, if $\langle c_{n}\rangle_{n}$ is one of those sequences, then that sequence will be simply referred as $c$. We will follow a similar rule for different sequences.
Lemma 5.7.
For every sequence of functionals $\langle c_{n}\rangle_{n}\preceq\langle e_{n}\rangle_{n}$ and a finite-to-one function $f$, there is a sequence of functionals $\langle d_{n}\rangle_{n}\preceq\langle c_{n}\rangle_{n}$ and a finite-to-one function $g$ such that
$(\forall{x\in\ell^{\infty}})\;|d_{n}(x)-d_{n}(x\circ g\circ f)|<2^{-n}.$
Proof.
Let $c^{\prime}$ and $g$ be obtained from Lemma 5.6. We are going to construct a sequence $\langle c^{m}\rangle$. Put $c^{0}_{m}=c_{m}$ for each $m$. Now let $c^{n+1}_{m}(x)=c^{n}_{m}(c^{\prime}(x))$. As $c^{n}\preceq e$, we have also $c^{n+1}\preceq c^{\prime}\preceq e$. In this way, and by Lemma 5.6, we ensure that $c^{n+1}_{m}(x\circ g\circ f)=c^{n}_{m}(x)$.
Now let
$d_{n}=\frac{1}{2^{n+2}}\big{(}c_{n}^{0}+c_{n}^{1}+c_{n}^{2}+\cdots+c_{n}^{2^{n+2}-1}\big{)}.$
Pick any $x\in\ell^{\infty}$. Notice that
$d_{n}(x\circ g\circ f)=\frac{1}{2^{n+2}}\big{(}c_{n}^{0}(x\circ g\circ f)+c_{n}^{0}(x)+c_{n}^{1}(x)+\cdots+c_{n}^{2^{n+2}-2}(x)\big{)}$
and so $|d_{n}(x)-d_{n}(x\circ g\circ f)|<2^{-n-1}+2^{-n-1}$. ∎
Theorem 5.8.
Assume $CH$. For any sequence of functionals $\langle c_{n}\rangle_{n}\prec\langle e_{n}\rangle_{n}$ there is a functional $\varphi$ in the intersection of $\{\overline{conv(c_{n}:n>k)}:k\in\omega\}$$\overline{\,\cdot\,}$ denotes a closure in the topology of pointwise convergence such that for any finite-to-one $f$ there is a finite-to-one $g$ with
$(\forall x\in\ell^{\infty})\;\;\varphi(x\circ g\circ f)=\varphi(x).$
Proof.
Let $\langle x_{\alpha}:\alpha\in\omega_{1}\rangle,\langle f_{\alpha}:\alpha\in\omega_{1}\rangle$ enumerate $\ell^{\infty}$ and all finite-to-one functions in $\omega^{\omega}$.
We construct a sequence $\langle c^{\alpha}\rangle_{\alpha\in\omega_{1}}$ with
  1. (1)
    $c^{\alpha}\preceq c$ for any $\alpha<\omega_{1}$,
  2. (2)
    $c^{\alpha}\preceq^{*}c^{\beta}$ for $\alpha>\beta$,
  3. (3)
    $\lim_{n}c_{n}^{\alpha}(x_{\alpha})$ exists,
  4. (4)
    there is a finite-to-one $g_{\alpha}$ such that
    $(\forall x\in l_{\infty})\;\;|c_{n}^{\alpha}(x\circ g_{\alpha}\circ f)-c_{n}^{\alpha}(x)|<\frac{1}{2^{n}}.$
First, for each $\alpha$, we define a sequence $\langle v^{\alpha}_{n}\rangle_{n}$ in the following way: Let $v_{n}^{0}=c_{n}$ and let $v_{n}^{\alpha}=c_{n}^{\beta}$ if $\alpha=\beta+1$. If $\alpha$ is limit, pick $\langle\alpha_{n}\rangle_{n}$ a sequence of ordinals increasing to $\alpha$. Assuming we have defined each $\langle c^{\xi}_{n}\rangle_{n}$ for each $\xi<\alpha$ we put
$v^{\alpha}=c^{\alpha_{n}}_{n}.$
Notice, that in each case $v^{\alpha}\preceq c$ and for any $\beta<\alpha$ we have $v^{\alpha}\preceq^{*}c^{\beta}$. Now, pick a subsequence $k_{n}$ such that $\lim_{n}v^{\alpha}_{k_{n}}(x_{\alpha})$ exists (there is one as $\langle v_{n}^{\alpha}(x_{\alpha})\rangle_{n}$ is a bounded sequence). Then, we apply Lemma 5.7 to $f_{\alpha}$ and the sequence $\langle v^{\alpha}_{k_{n}}\rangle_{n}$ to obtain $g_{\alpha}$ and $\langle c_{n}^{\alpha}\rangle_{n}$. This finishes our recursive construction.
Finally, let $\varphi(x)=\lim_{n}c_{n}^{\alpha}(x)$ for the minimal $\alpha<\omega_{1}$ for which this limit exists. According to Fact 2.10, for any $\beta$ above such $\alpha$, we have $\varphi(x)=\lim_{n}c_{n}^{\beta}(x)$. Thanks to (3) we know that $\varphi$ is defined on each element of $\ell^{\infty}$. A simple reflection argument shows that $\varphi$ is indeed a functional.
We will show that $\varphi\in\bigcap\{\overline{conv(c_{n}:n>k)}:k\in\omega\}$: Pick any $x\in\ell^{\infty}$. There is $\alpha<\omega_{1}$ s.t. $x=x_{\alpha}$. Thus $\varphi(x)=\lim_{n}c_{n}^{\alpha}(x)$ and so for any $k\in\omega$, we have that $\varphi(x)\in\overline{conv(c_{n}^{\alpha}(x):n>k)}$. Hence $\varphi\in\overline{conv(c_{n}^{\alpha}:n>k)}\subseteq\overline{conv(c_{n}:n>k)}$.
To see the last part, let $f$ be a finite-to-one function. There is $\alpha$ such that $f_{\alpha}=f$. Pick any $k\in\omega$. For every functional $\psi\in conv(c^{\alpha}_{n}:n>k)$ we have
$|\psi(x\circ g_{\alpha}\circ f_{\alpha})-\psi(x)|\leqslant\frac{1}{2^{k}}.$
So the same is true for $\varphi$ since $\varphi\in\overline{conv(c^{\alpha}_{n}:n>k)}$. ∎
As a corollary we get Theorem 5.5 promised at the beginning of this section:
Corollary 5.9.
There is an RB-minimal shift-invariant measure extending the density.
Now, we will show that an simple modification of the argument above yields a similar result for the Rudin-Keisler order; the main idea is to add a restriction to additionally ask that the resulting measure is a P-measure.
In what follows, for $c$ being a functional on $\ell^{\infty}$ and $A\subseteq\omega$, by $c(A)$ we mean $c(\chi_{A})$.
Lemma 5.10.
Suppose $(c_{n})\preceq(e_{n})$ and let $(A_{k})$ be a partition of $\omega$ such that $r_{k}\colon=\lim_{n}c_{n}(A_{k})$ exists for every $k$. Then there is a sequence $(n_{i})$ and $P\subseteq\omega$ such that
  • (1)
    $A_{k}\subseteq^{*}P$ for each $k$,
  • (2)
    $\lim_{i}c_{n_{i}}(P)=\sum_{k}r_{k}$.
Proof.
Using the assumptions we can easily find an increasing sequence $(n_{i})$ such that
$|c_{n_{i}}(A_{0}\cup\dots\cup A_{i})-(r_{0}+\dots+r_{i})|<1/2^{i}.$
Let
$F_{i}=\mathrm{supp}(c_{n_{i}})\setminus(A_{0}\cup\dots\cup A_{i})$
and finally let $P=\omega\setminus\bigcup_{i}F_{i}$.
Clearly, (1) holds.
We check (2). Let $\varepsilon>0$. Find $j$ big enough so that
  • 1/$21/2^{j}<\varepsilon/$ 2,
  • $\sum_{k>j}r_{k}<\varepsilon/2$.
If $i>j$, then
$\Big{|}c_{n_{i}}(A_{0}\cup\dots\cup A_{i})-\sum_{k}r_{k}\Big{|}\leq\Big{|}c_{n_{i}}(A_{0}\cup\dots\cup A_{i})-\sum_{k\leq i}r_{k}\Big{|}+\Big{|}\sum_{k}r_{k}-\sum_{k\leq i}r_{k}\Big{|}<\varepsilon.$
But
$c_{n_{i}}(P)=c_{n_{i}}(A_{0}\cup\dots\cup A_{i}),$
and so $|c_{n_{i}}(P)-\sum_{k}r_{k}|<\varepsilon$. ∎
Theorem 5.11.
(CH) There is a shift-invariant RK-minimal measure.
Proof.
We proceed as in the proof of the RB-theorem but additionally we enumerate all the infinite partitions $\{(A^{\alpha}_{k})\colon\alpha<\omega_{1}\}$ of $\omega$ in the way that each partition appears cofinally many times.
At step $\alpha$, having defined $(c^{\alpha}_{n})$ we add the following step. If for every $k$ the limit $\lim_{n}c^{\alpha}_{n}(A_{k})$ exists, then we use the above lemma and we pass to an appropriate subsequence of $(c^{\alpha})_{n}$. Otherwise, we do not change the sequence $(c^{\alpha})_{n}$.
In this way we finally obtain a (shift-invariant, RB-minimal) P-measure: for each infinite partition $(A_{k})$ of $\omega$ we will find $\alpha$ such that $(A_{k})=(A^{\alpha}_{k})$ and $\lim_{n}c^{\alpha}(A_{k})$ does exist for each $k$. Then $P$ promised by the lemma works.
But a measure $\nu$ which is RB-minimal and P is in fact RK-minimal. Indeed, if $\mu\leq_{RK}\nu$ then $\mu\leq_{RB}\nu$ (because $\nu$ is a P-measure, see [BNCMM25, Fact 2.6]). Then, by RB-minimality, we have $\nu\leq_{RB}\mu$, and thus also $\nu\leq_{RK}\mu$. ∎
Remark 5.12.
Let $\mu$ be a RB-minimal shift-invariant measure. Let $f$ be finite-to-one. If the lengths of intervals contained in $f^{-1}[\{n\}]$ are not bounded, then by argument as in Proposition 5.4 we have that there is no $\mu$-positive selector for $\langle f^{-1}[\{n\}]\rangle$. In particular if $gf[\mu]=\mu$, then $\mu(\mathrm{Fix}(gf))=0$. Thus, there is a $\mu$-positive set $A$ with $gf[A]\cap A=\emptyset$. However $gf$ in some sense preserves the structure of $\mu$ on subsets of $A$.
What this shows is that while Q-measures obtain RB-minimality by ignoring a lot of the domain (having ’rare’ sets of measure 1), this $\mu$ achieves the same by having a lot of nontrivial inner ’symmetries’.
We finish this section with several remarks concerning the measurability of the measures.
Remark 5.13.
Modifying the above construction we may additionally assume that the functional promised by Theorem 5.8 is universally measurable. Indeed, begin with enumerating all bounded Radon measures on $2^{\omega}$ as $\langle\mu_{\alpha}:\alpha<\omega_{1}\rangle$. Then given $\langle v_{n}^{\alpha}\rangle$ we pick a subsequence ensuring that $\lim_{n}v_{k_{n}}^{\alpha}(x)$ exists $\mu_{\alpha}$-almost everywhere. This can be done by the reflexivity of the space $L^{2}(l_{\infty},\mu_{\alpha})$ in which $\langle v_{n}^{\alpha}\rangle_{n}$ is a bounded sequence. Then, using Mazur’s lemma, we may obtain that this sequence converges in norm. More details on this approach can be found in [God16].
This means that measurability is yet another property distinguishing Q-measures with the measure of Theorem 5.5. We enclose a sketch of the proof of the following proposition although it follows from much more general result of Mokobodzki who proved that rapid filters are not Lebesgue measurable (see [Bar92]).
Proposition 5.14.
If $\mathcal{F}$ is a Q-filter, then it is not Lebesgue measurable. Consquently, Q-measures are not Lebesgue measurable.
Proof.
Since the only possiblity for a filter to be measurable is to have Lebesgue measure 0, it is enough to show that $\mathcal{F}$ intersects every closed subset of $2^{\omega}$ of positive Lebesgue measure. In other words, we have to show that if $U$ is an open set which is not of full measure, then there is $X\in\mathcal{F}\setminus U$. Let $U=\bigcup_{n}C_{n}$, where each $C_{n}$ is a basic clopen.
Without loss of generality we may assume that that for each $n$ there is a finite $F_{n}$ such that $C_{n}=\{x\in 2^{\omega}\colon\forall i\in F_{n}\ x(i)=1\}$. Also, we may assume that $(F_{n})$ are pairwise disjoint and of cardinality at least $2$. Applying Q-property, we get that there is $X\in\mathcal{F}$ such that $|X\cap F_{n}|\leq 1$ for every $n$. Clearly, $X\notin U$. ∎
Corollary 5.15.
No universally measurable measure has a Q-measure RK below.
Proof.
We just have to see that if $\mu$ is universally measurable and $\nu\leq_{RK}\mu$, then $\nu$ is universally measurable, too. Then the corollary follows from Proposition 5.14.
So, let $\mu$ be universally measurable, and let $\nu=f[\mu]$ for some function $f\colon\omega\to\omega$. Notice that the function $h\colon 2^{\omega}\to 2^{\omega}$ defined by $h(\chi_{A})=\chi_{f^{-1}[A]}$ is continuous and so it is universally measurable. So, the measure $\nu=\mu\circ h$ is a composition of universally measurable functions, and so it is universally measurable (see e.g. [Fre06, 434Df]). ∎
One can ask how badly measurable are Q-filters (e.g. what about measurability with respect to other reasonable measures on the Cantor set). In the light of Proposition 5.14 it may be surprising that the selective ultrafilters seem to be easier to measure with respect to some measures than other ultrafilters:
Remark 5.16.
Bartoszynski in [Bar92] considered measures on $2^{\omega}$ of the form $\mu_{p}=\prod_{n}p_{n}\delta_{0}+(1-p_{n})\delta_{1}$ (notice that if $p$ is constantly $1/2$, then $\mu_{p}$ is the standard Haar measure on $2^{\omega}$). It turns out that if $(p_{n})$ converges quickly to 00, then all the filters are $\mu_{p}$-measurable. In particular, Bartoszynski showed that if $\sum_{n}p_{n}<\infty$, then all the filters are $\mu_{p}$-measurable. It is not clear to us if there is a sequence $(p_{n})$ converging to 00 such that some ultrafilter $\mathcal{U}$ is not $\mu_{p}$-measurable. But it is not hard to see that selective ultrafilters are definitely $\mu_{p}$-measurable for every such sequence. Indeed, by a result of Mathias (see [Mat77]) a selective ultrafilter intersects every tall analytic ideal. So, if $\mathcal{U}$ is selective, then there is $X\in\mathcal{U}$ such that $\sum_{n\in X}p_{n}<\infty$. By [Bar92, Theorem 1.3] it means that $\mathcal{U}$ is $\mu_{p}$-measurable.